The ratio of Mike's age to his brother Sam's age is 2:3. In 10 years the ratio will be 4:5. How old is Mike?
Answer by HasanSahin(52)

Let's say Mike's age is now 2x and Sam's age is 3x.
After 10 years Mike will be 2x + 10 and Sam will be 3x + 10
The ratio will be 4:5 and if you multiply Mike's age with 5 and Sam's age with 4 you'll equal them such that >> (2x + 10)*5 = (3x + 10)*4
Therefore, 10x + 50 = 12x + 40 >> 2x = 10 >> x=5
Let's find the age of Mike >> 2x = 10 years old..
RF.